40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 1197     Entry time: Sat Feb 7 03:50:14 2015
Author: Zach 
Type: DailyProgress 
Category: SiFi - ringdown 
Subject: Electrical connections fixed, clamp adjusted, chamber repumped and cooling 

I vented the chamber today to redo the clamping and investigate our wiring issues.

Clamp

Since I observed relatively low Q even at cryogenic temperatures, I assumed there was some jankiness with how we clamped the cantilever when we installed the sapphire washers. Recall that the lower part of the steel clamp has a circular lip near the center around the screw hole, and it was too wide to allow the sapphire washers to fit around it. This meant that the lower washer was only being held by this lip, and not by the full surface area of the clamp. Also, when we installed the washers, we didn't remove the smallest can around the physics package, so we were doing a bit of guesswork as to how well aligned the entire clamp stack was. This meant that there could have been some slight rubbing, for example. Here is a photo of what it looked like in profile when I did remove the can today:

You can see what I mean about the lip, and it's also clear that the stack was not very well aligned. To fix the lip problem, I found a steel washer that was just about the right thickness and drilled the center hole out wide enough that it fit around the lip. This way, the lower sapphire washer will be supported by a larger surface from below (of course, the real solution will be to either design a new clamp or get wide-enough-ID sapphire washers). The picture on the left below shows the washer around the lip.

There was also some dust and other gunk visible to the eye, so I thoroughly cleaned all parts in the stack with methanol and isopropanol. I then carefully restacked the components and reclamped (a little tighter than we did last time, as well). The final stack is shown below at right.

 

Wiring

I checked each connection from the feedthrough to the heater or RTD, and found that everything seemed to be in order, so there must have just been a short when we closed up last time. I wrapped some extra kapton around each connector solder joint to provide insulation and extra strain relief, and everything stayed as it should be when I resealed the chamber. I *did* accidentally break a joint on the wire for the ESD while closing up---whoops---but I decided it was more hassle to fix it than necessary for this next run. I'll resolder it when we cycle again.

 

The chamber is under vacuum now and I filled the reservoir with nitrogen. The clamp was at 200 K when I left around 10pm, so I'm hoping things will be calm and cool when I come in tomorrow.

 

ELOG V3.1.3-